OpenStudy (anonymous):

Show the work to solve |5x – 2| greater than or equal to 8

OpenStudy (anonymous):

5x - 2 > 8 or 5x -2<-8\[x>2 or x < -6/5\]

OpenStudy (anonymous):

\[(-\infty,\frac{-6}{5}]U[2.\infty)\]

OpenStudy (anonymous):

\cup

OpenStudy (anonymous):

\[(-\infty, -\frac{6}{5}]\cup [2,\infty)\]

OpenStudy (anonymous):

\[\Huge{\cup}\]\[Thanks : )\]

OpenStudy (anonymous):

as opposed to \cap

OpenStudy (anonymous):

\[A\cap B\]